Circular No. 7431

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000ck IN IC 4355 T. Puckett, Mountain Town, GA, reports the discovery of an appar-

ent supernova (mag 17.0) on an unfiltered CCD frame (limiting mag 20.2) taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on May 26.28 UT. SN 2000ck is located at α = 13^{h}58^{m}06^{s}.15, δ = +28^{o}25′13′′.0 (equinox 2000.0), which is 0′′.3 west and 8′′.8 south of the center of IC 4355. The new object was also present on an unfiltered CCD frame taken on May 27.09, but it does not appear on Palomar Sky Survey images taken on 1996 July 9 (limiting mag ∼ 21.0) or 1988 Mar. 20 (limiting mag ∼ 21.0).

COMET C/2000 K1 (LINEAR) Additional observations on May 27 by L. Amburgey (Fitchburg, MA)

and F. B. Zoltowski (Edgewood, NM), together with May 3 prediscovery astrometry from LINEAR and the following parabolic orbital elements, appear on MPEC 2000-K26.

T = 2000 Jan. 27.200 TT  ω = 19.^{o}789
Ω = 260.360 }2000.0
q = 6.32915 AU  i = 117.032
2000TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
May 26 16^{h}36.^{m}26 + 9^{o}41.^{′}3 5.495 6.382 148.^{o}6 4.^{o}7 15.7
June 5 16 27.72 +10 50.1 5.524 6.391 146.0 5.1 15.8
15 16 19.39 +11 47.2 5.586 6.401 140.4 5.8 15.8
25 16 11.58 +12 32.3 5.676 6.411 132.9 6.7 15.8

COMET C/2000 K2 (LINEAR) Additional observations on May 27 by Zoltowski, together with May 4

prediscovery astrometry from LINEAR and the following parabolic orbital elements, appear on MPEC 2000-K27.

T = 2000 Oct. 12.024 TT  ω = 107.^{o}183
Ω = 195.247 }2000.0
q = 2.43102 AU  i = 25.634
2000 May 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green